OpenStudy (anonymous):

Find the value of x given a set of numbers is 12, 56, 3, 9, 18, 21, x, 45 and the average is 32.

OpenStudy (tkhunny):

How does one calculate an average of 8 values?

OpenStudy (anonymous):

by adding then dividing

OpenStudy (tkhunny):

What's stopping you from doing that with your 8 values?

OpenStudy (anonymous):

not knowing what one value is

OpenStudy (tkhunny):

Fair enough. Just ignore that for now and proceed, anyway. See if anything useful happens. You already know what the average is. Maybe that will help.

OpenStudy (anonymous):

okay so i know the avergge is 32 but how is that any help?

OpenStudy (tkhunny):

You're not going to do it, are you... \(\dfrac{12+56+3+9+18+21+x+45}{8} = 32\) Are you SURE that's not helpful?

OpenStudy (dwelsher):

(12+56+3+9+18+21+x+45)/8=32 164+x=256 x=256-164 x=92

OpenStudy (anonymous):

No, I am going to do it I'm just really confused in this. After I added it I get \[164+x \div8=32\]

OpenStudy (tkhunny):

Add the parentheses around (164 + x) then multiply everything by 8. A little faith got us there. :-)

OpenStudy (anonymous):

@dwelsher how did you get 256 on the second line?

OpenStudy (tkhunny):

Really? What's 8 * 32? You MUST pay better attention.